# Atomic clock interferometry using optical tweezers

### v1

@@ Abstract
Clock interferometry refers to the coherent splitting of a clock into two different paths and recombining in a way that reveals the proper time difference between them. In contrast to comparing two separate clocks, this type of measurement can test quantum gravity theories. Atomic clocks are currently the most accurate time keeping devices. Here we propose using optical tweezers to implement clock interferometry. Our proposed clock interferometer employs an alkaline-earth-like atom held in an optical trap at the magic wavelength. Through a combination of adiabatic, tweezer-based, splitting and recombining schemes and a modified Ramsey sequence on the clock states, we achieve a linear sensitivity to the gravitational time dilation. Moreover, the measurement of the time dilation is insensitive to relative fluctuations in the intensity of the tweezer beams. We analyze the tweezer clock interferometer and show that it is feasible with current technological capabilities. The proposed interferometer could test the effect of gravitational redshift on quantum coherence, and implement the quantum twin paradox.
@@ I Introduction
Atom Interferometers (AIFs) are techniques for coherently manipulating spatial superpositions of atomic wavefunctions and measuring their phase difference. AIFs come in many forms, yet they all share the same basic steps and structure [1]. The AIF consists of input and output ports, typically two of each type. The ports may be characterized by the outcome of different observables, such as position [2], momentum [3], and angular momentum [4]. A single atomic wavepacket populating one of the input ports is coherently split into two spatially separated wavepackets, so that each wavepacket is traveling on a different path. Each path may be under the influence of a different potential, so that each wavepacket acquires a different phase. After acquiring the phase, the wavepackets are coherently recombined. This process maps the phase difference between the wavepackets to population probabilities at the output ports. The populations are, in turn, measured to infer the phase difference.
@@
The most common type of AIFs is free space light-pulse interferometry, such as the Kasevich-Chu Interferometer (KCI) [3]. In such architectures, atoms in free fall are split and recombined in momentum space through the absorption of photons, thus creating two possible spatial trajectories. An alternative approach is to confine the wave packets in a potential well during all or part of their trajectories. These types of AIFs, also known as guided AIFs, offer several potential advantages over traditional light-pulse AIFs, including arbitrary atom trajectories, precise positioning, long probing times, and compact experimental setups. In recent years, several architectures for guided AIFs have been proposed. Among the methods available for manipulating atomic trajectories are optical lattices [5] and optical tweezers [2, 6, 4].
@@
AIFs, particularly KCI, have been successfully employed in numerous precise measurements [7, 8]. The fact that AIFs involve massive particles and can detect extremely small variations in force fields makes it an excellent choice for gravitational measurements. Indeed, AIFs have been used to test the weak equivalence principle [9] and to determine the gravitational constant [10, 11]. One central tenet of general relativity is the modification of proper time due to masses. Proper time, denoted by $\tau$, is the time measured by an ideal clock moving with the reference frame. The theory of General Relativity predicts that clocks tick slower near a large mass, a prediction confirmed in several experiments comparing independent clocks positioned at different distances relative to Earth’s mass [12, 13, 14, 15, 16, 17].
@@
In principle, time dilation can also be probed using AIF. As proper time is inherently tied to the metric in general relativity, observing its influence on the wavefunction’s interference pattern could probe the regime in which both general relativity and quantum mechanics have a measurable effect [18], allowing tests of quantum gravity theories. As an example, some theories propose treating proper time as a quantum operator, with mass as its conjugate [19, 20]. In these theories, the uncertainty in proper time is predicted to affect the visibility of the interference pattern [21]. Other proposals suggest interferometric measurements to test the connection between gravity and decoherence on macroscopic scales [22, 23].
@@
where $m$ is the mass, $c$ is the speed of light, and $\hbar$ is the reduced Planck constant. This relation may lead one to conclude that it is possible to use the accumulated phase in AIF as a means to measure the difference of the proper time. This claim was made by Muller et al. as they reinterpreted light-pulse interferometry experiments as a clock ticking in the Compton frequency, and as such, used its measured phase difference for calculating the gravitational red shift [25]. However, it was claimed that the proper time cannot be measured by a particle with a single internal state [26]. The phase difference in a light-pulse AIF in free fall is only due to the different phases of light pulses relative to the atoms accelerated by gravity, $g$ [24]. To ascertain the proper time difference, it is necessary to employ a clock, which requires at least two internal states. Atomic clock interferometry (ACIF) has not yet been performed.
@@
A step towards ACIF has been taken using the ${}^{1}S_{0}-{}^{3}P_{0}$ optical clock transition of ${}^{88}$r [30, 31, 32]. In these experiments, however, the splitting was accomplished using a single photon optical transition. Consequently, the atoms in each interferometer arm were in one of the clock states and not in a superposition of them. Therefore, they did not act as a clock in the sense required for measuring the proper time difference between the arms. Guided ACIFs hold a great promise for gravitational measurements, in particular for testing the Universality of the Gravitational Redshift (UGR) – the principle that the gravitational time dilation is independent of the inner workings of the clock [33, 34]. In most light pulse interferometers, the proper time difference between the arms is zero [24, 35, 36]. Previous works were focused mainly on suggesting ACIFs based on light-pulse schemes that overcome this difficulty [34, 33, 37]. Guided interferometers, on the other hand, are inherently well-suited for measuring differences in proper time. [34]. However, despite the strong motivation to test theories of quantum gravity, a concrete scheme for a guided ACIF has yet to be proposed.
@@ II Tweezer atomic interferometry
In Ref. [2], we introduced a guided AIF scheme using optical tweezers. As the ACIF scheme we present in section III builds upon this concept, we briefly revisit it here for clarity. In the tweezer AIF, the atom’s motion is controlled throughout the entire interferometric sequence by manipulating the position of the optical tweezers. The sequence starts with the atoms located in one of the tweezers. Coherent splitting and recombining are achieved by an adiabatic change of the traps’ positions. We proposed two configurations for the splitters and combiners, utilizing either a two-tweezer or a three-tweezer setup.
@@
where $\sigma$ is the waist of the Gaussian beam creating each of the tweezers, $V_{0}$ is the potential depth, and $\Delta(t)$ is the depth difference (detuning) between the two traps. Splitting of the wavefunction is achieved by changing the separation between the traps, $d(t)$, and the detuning in the following manner:
@@
The three-tweezer scheme can be explained using the same principle of adiabatic following of the atomic wavefunction. The initial potential has three traps: two at distance $d_{max}$ from both sides of a central trap, and with a detuning $\Delta$ (the central trap is deeper). The atom is initially placed at the center trap, in the ground state of the triple-trap potential. The splitting follows a similar two steps process. First, the distance of the external traps from the center trap is decreased to $d_{min}$ while the detuning is lowered to zero. Then, the external traps are distanced to $d_{max}$ while the detuning is changed to $-\Delta$. The atom ends at the ground state of the final potential, which is the symmetric superposition of being in the two external traps.
Similarly to the two-tweezer case, the combiner is the time-reversal of the splitter sequence. Due to the symmetry of the Hamiltonian, the combiner maps only the symmetric part of the final wavefunction back to the central trap, and leaves the anti-symmetric part unchanged. Therefore, the symmetric part of the final wavefunction is mapped to the central trap and the anti-symmetric part is mapped to the external traps. Any phase difference between the arms maps to a superposition of being in the central and external traps. The larger number of degrees of freedom in the three-tweezer splitter and combiner provides the important benefit of error detection, that is, the possibility to detect and discard runs where, due to experimental imperfections, the splitting or recombining processes were not successful.
@@
The guided interferometer architecture with optical tweezers offers several advantages over free-fall light-pulse interferometry. One of the biggest limitations of light-pulse interferometry is that since the atoms are in free-fall and the acquired phase scales as $T^{2}$ [39], better sensitivity requires increasingly larger apparatus. The guided interferometer allows for a long interrogation time while remaining compact. Another limitation of light-pulse interferometry is that the atoms are in free-fall. In the proposed tweezer AIF, the atom’s position is entirely determined by the tweezer position, which allows for arbitrary trajectories during phase accumulation. Due to its enhanced sensitivity and arbitrary atom trajectory, the tweezer AIF can measure effects in classical and quantum gravity, e.g measuring the gravitational constant $G$ [40], testing the Newtonian law at small distances [40], measuring the gravitational redshift [34] and searching for evidence of the gravitational-field quantization [41, 42]. Most importantly, in the context of this paper, guided interferometers are sensitive to proper time differences between the arms [34], while existing light-pulse interferometers are inherently insensitive to proper time [36]. Moreover, an optical tweezer AIF working at a magic wavelength with an atomic clock is ideally suited to measure the gravitational time dilation, as we demonstrate in section III.
@@ III Tweezer clock interferometry
Optical tweezer AIFs, like those discussed in section II, can be used to measure phases arising from proper time differences if the atom has at least two internal states. This is due to the fact that an atom with a single internal degree of freedom cannot produce a periodic signal necessary for measuring proper time [26, 43]. Interferometers employing only a single internal state can only measure effects stemming from a difference of the gravitational potential at different location, analogous to a gravitational Aharonov-Bohm effect [44]. Such observations were made using atomic and neutron interferometry [45]. Atomic clock interferometry was previously proposed as a method to measure proper time differences [33, 37, 34, 43]. However, these proposals suffer from the same limitation of free falling AIF discussed in section II. In what follows, we present a proposal for a guided ACIF, using optical tweezers at a magic wavelength. Our scheme inherits the advantages of the tweezer AIF, namely long probing duration and high position accuracy. Additionally, the measurement of the proper time difference is insensitive to relative intensity fluctuations, which makes it applicable to current tweezer technology.
@@
Our guided ACIF scheme is depicted in Fig. 1. It is a modification of the scheme discussed in section II, and can be applied with both the two- and three-tweezer approaches. It employs an atom which can be used as an atomic clock, i.e., with two long-lived internal states, denoted by |$\ket{g}$ and |$\ket{e}$. We choose the tweezers to operate at the magic wavelength of the clock, where both clock states experience the same trap potential. The atom is initially prepared in the ground state |$\ket{g}$. Before the wavepacket splitting stage, a $\pi/2$-pulse of an optical field, which is resonant with the |$\ket{g}\leftrightarrow\ket{e}$ transition, is applied. This pulse generates the superposition $\frac{1}{\sqrt{2}}(\ket{g}+\ket{e})$.
Following the $\pi/2$-pulse, the spatial splitting of the clock wavefunction takes place, using either two or three tweezers. Each interferometer arm is moved to a different position in the gravitational field. Therefore, each part of the clock wavefunction experiences a different proper time. Before the recombining phase, a $\pi$-pulse on the |$\ket{g}$-|$\ket{e}$ transition is applied to only one of the arms. The sequence ends with the spatial recombining of the wavefunction, as described in section II, followed by another $\pi/2$-pulse on the clock degrees of freedom to close the Ramsey sequence. The measured observables are the internal state of the atom and the spatial output port. As we demonstrate below, both quantities are required for a measurement of proper time differences in a coherent superposition of an atom.
@@
After the splitting and before the recombining, we define the energy difference $\Delta E$ (and the corresponding frequency $\omega_{0}$) between the two clock states in path 2 as
$\Delta E=h\omega_{0}=E_{e;2}-E_{g;2}\,.$ (7)
@@
$E_{e;1}-E_{g;1}=\Delta E+h\epsilon$ (8)
with $\epsilon=\frac{\omega_{0}}{c^{2}}\Delta\varphi$ being the frequency difference due to the gravitational redshift. Our goal for the ACIF sequence is to allow determination of $\epsilon$. It is important to note that in the way we wrote Eqs.(7-8) we explicitly assumed that the energy difference between the clock states does not depend on the depth of the confining optical potential, which is justified when the trap is operated at the magic wavelength.
@@
$U_{\text{phase}}(T)=\begin{pmatrix}e^{-\frac{i}{\hbar}E_{g;1}T}&0&0&0\\ 0&e^{-\frac{i}{\hbar}E_{g;2}T}&0&0\\ 0&0&e^{-\frac{i}{\hbar}E_{e;1}T}&0\\ 0&0&0&e^{-\frac{i}{\hbar}E_{e;2}T}\end{pmatrix}\,.$ (12)
@@
|$\ket{\psi_{f}}=U_{\pi/2}^{\dagger}U_{BS}^{\dagger}U_{\pi,l}U_{\text{phase}}(T)U_{BS}U_{\pi/2}\ket{g;1}\,.$ (13)
Following the sequence there are two types of observables one can measure. The first observable is the spatial output port of the atom. The probability of the atoms to exit the interferometer from the upper port is given by
@@
$\begin{split}&P_{1}(\psi_{f})=P_{g;1}+P_{e;1}=|\langle g;1|\psi_{f}\rangle|^{2}+|\langle e;1|\psi_{f}\rangle|^{2}=\\ &=\frac{1}{2}\left[1-\sin\left(T\left(E_{g;1}-E_{g;2}+\frac{\epsilon}{2}\right)\right)\sin\left(T\left(\omega_{0}+\frac{\epsilon}{2}\right)\right)\right]\,.\end{split}$ (14)
@@
A crucial requirement for an interferometric measurement is the ability to verify that the atom was indeed in a coherent superposition during the sequence. The result of Eq.(14) allows this, since it exhibits coherent oscillations between the exit ports as a function of $T$. In contrast, if the wave packet collapses randomly to one of the paths and the state becomes completely mixed, the exit port is $P_{1,\text{mixed}}(\psi_{f})=\frac{1}{2}$. This means that measuring oscillations around the probability of $\frac{1}{2}$ as a function of time verifies the coherence of the atomic wavefunction.
The main disadvantage of measuring the spatial output port is the exit probability dependence on the difference between the eigenenergies: $E_{g;1}-E_{g;2}$. This makes this observable sensitive to relative fluctuations between the intensity of the trap beams. There are two ways to cope with this issue. On one hand, one can set stringent requirements on the laser intensity stability throughout the interferometer time which will ensure low noise in this measurement. This approach can be very challenging experimentally. An alternative approach is to acknowledge that the fluctuations exist and result in almost a random phase in the first sine term of Eq.(14). The probability distribution, $P(x)$, of $x=\sin(\phi)$ with $\phi$ being evenly and randomly distributed is $P(x)=(\pi\sqrt{1-x^{2}})^{-1}$. Thus, if the wave packet is coherently split and held, the distribution of $P_{1}$ should follow $P(x)$ over a dataset large enough.
The second available observable is the inner clock state of the atom. The probability of the atoms to be in the clock ground state at the end of the sequence is given by
@@
$\begin{split}P_{g}(\psi_{f})&=P_{g;1}+P_{g;2}=|\langle g;1|\psi_{f}\rangle|^{2}+|\langle g;2|\psi_{f}\rangle|^{2}\\ &=\frac{1}{2}\left[1-\sin\left(T\left(\omega_{0}+\frac{\epsilon}{2}\right)\right)\sin\left(T\frac{\epsilon}{2}\right)\right]\,.\end{split}$ (15)
@@
As is readily seen from Eq.(15), in contrast to the case of measuring the spatial output port, the probability to finish in the ground state depends only on $\omega_{0}$ and $\epsilon$, but not on the eigenenergies separately. This means that this observable is insensitive to relative fluctuations in the depths of the tweezers, which would shift the eigenenergies, but do not affect $\omega_{0}$, as long as we operate at the magic wavelength.
@@
The redshift is extracted from Eq.(15). By scanning the waiting time, $T$, the probability to find the atoms in the ground state oscillates due to interference of the two paths. The oscillation frequency of this interference pattern is the average clock frequency in the two paths $\omega_{0}+\frac{\epsilon}{2}$. Its visibility is given by

### v3

@@ Abstract
Clock interferometry refers to the coherent splitting of a clock into two different paths and recombining in a way that reveals the proper time difference between them. Unlike the comparison of two separate clocks, this approach allows testing how non-flat spacetime influences quantum coherence. Atomic clocks are currently the most accurate time keeping devices. Here we propose using optical tweezers to implement clock interferometry. Our proposed clock interferometer employs an alkaline-earth-like atom held in an optical trap at the magic wavelength. Through a combination of adiabatic, tweezer-based, splitting and recombining schemes and a modified Ramsey sequence on the clock states, we achieve a linear sensitivity to the gravitational time dilation. Moreover, the measurement of the time dilation is insensitive to relative fluctuations in the intensity of the tweezer beams. We analyze the tweezer clock interferometer and show that it is feasible with current technological capabilities. The proposed interferometer could test the effect of gravitational redshift on quantum coherence, and implement the quantum twin paradox.
@@ I Introduction
Atom Interferometers (AIFs) are techniques for coherently manipulating spatial superpositions of atomic wavefunctions and measuring their phase difference. AIFs come in many forms, yet they all share the same basic steps and structure [1]. The AIF consists of input and output ports, typically two of each type. The ports may be characterized by the outcome of different observables, such as position [2], momentum [3, 4], and angular momentum [5]. A single atomic wavepacket populating one of the input ports is coherently split into two spatially separated wavepackets, so that each wavepacket is traveling on a different path. Each path may be under the influence of a different potential, so that each wavepacket acquires a different phase. After acquiring the phase, the wavepackets are coherently recombined. This process maps the phase difference between the wavepackets to population probabilities at the output ports. The populations are, in turn, measured to infer the phase difference.
@@
The most common type of AIFs is free space light-pulse interferometry, such as the Kasevich-Chu Interferometer (KCI) [6]. In such architectures, atoms in free fall are split and recombined in momentum space through the absorption of photons, thus creating two possible spatial trajectories. An alternative approach is to confine the wavepackets in a potential well during all or part of their trajectories. These types of AIFs, also known as guided AIFs, offer several potential advantages over traditional light-pulse AIFs, including arbitrary atom trajectories, precise positioning, long probing times, and compact experimental setups. An early demonstration of a guided AIF was conducted with a Bose-Einstein Condensate in a double-well potential [7]. In recent years, several alternative architectures for guided AIFs have been proposed. Among the proposed methods for guided atomic interferometry, some techniques manipulate atomic trajectories using optical lattices [8, 5], while other utilize optical tweezers [9, 2, 10]. A significant challenge in guided interferometry lies in fulfilling the stringent requirement for the intensity stability of the confining optical potential, which is crucial to prevent differential phase noise.
@@
AIFs, particularly KCI, have been successfully employed in numerous precise measurements [11, 12]. The fact that AIFs involve massive particles and can detect extremely small variations in force fields makes it an excellent choice for gravitational measurements. Indeed, AIFs have been used to test the weak equivalence principle [13] and to determine the gravitational constant [14, 15]. One central tenet of general relativity is the modification of proper time due to the metric. Proper time, denoted by $\tau$, is the time measured by an ideal clock moving with the reference frame. The theory of General Relativity predicts that clocks tick slower near a large mass, a prediction confirmed in several experiments comparing independent clocks positioned at different distances relative to Earth’s mass [16, 17, 18, 19, 20, 21].
@@
In principle, time dilation can also be probed using AIF. As proper time is inherently tied to the metric in general relativity, observing its influence on the wavefunction’s interference pattern could probe the regime in which both general relativity and quantum mechanics have a measurable effect [22]. As an example, some theories propose treating proper time as a quantum operator, with mass as its conjugate [23, 24]. In these theories, the uncertainty in proper time is predicted to affect the visibility of the interference pattern [25]. Other proposals suggest interferometric measurements to test the connection between gravity and decoherence on macroscopic scales [26, 27].
@@
where $m$ is the mass, $c$ is the speed of light, and $\hbar$ is the reduced Planck constant. This relation may lead one to conclude that it is possible to use the accumulated phase in AIF as a means to measure the difference of the proper time. This claim was made by Muller et al. as they reinterpreted light-pulse interferometry experiments as a clock ticking in the Compton frequency, and as such, used its measured phase difference for calculating the gravitational red shift [29]. However, it was claimed that the proper time cannot be measured by a particle with a single internal state [30]. The phase difference in a light-pulse AIF in free fall is only due to the different phases of light pulses relative to the atoms accelerated by gravity, $g$ [28]. To ascertain the proper time difference, it is necessary to employ a clock, which requires at least two internal states. While atomic clock interferometry (ACIF) has been previously realized with hyper-fine states [31], it has yet to be performed using states separated by optical transitions. Having a large transition frequency is a crucial requirement for a feasible measurement of the gravitational redshift.
@@
A step towards ACIF has been taken using the ${}^{1}S_{0}-{}^{3}P_{0}$ optical clock transition of 88Sr [35, 36, 37]. In these experiments, however, the splitting was accomplished using a single photon optical transition. Consequently, the atoms in each interferometer arm were in one of the clock states and not in a superposition of them. Therefore, they did not act as a clock in the sense required for measuring the proper time difference between the arms. Guided ACIFs hold a great promise for gravitational measurements, in particular for testing the Universality of the Gravitational Redshift (UGR) – the principle that the gravitational time dilation is independent of the inner workings of the clock [38, 39]. In most light pulse interferometers, the proper time difference between the arms is zero [28, 40, 41]. Previous works were focused mainly on suggesting ACIFs based on light-pulse schemes that overcome this difficulty [39, 38, 42]. Guided interferometers, on the other hand, are inherently well-suited for measuring differences in proper time. [39]. However, a concrete scheme for a guided ACIF using optical transitions has yet to be proposed.
@@ II Tweezer atomic interferometry
In Ref. [2], we introduced a guided AIF scheme using optical tweezers. As the ACIF scheme we present in section III builds upon this concept, we briefly revisit it here for clarity. In the tweezer AIF, the atom’s motion is controlled throughout the entire interferometric sequence by manipulating the position of the optical tweezers [43]. The sequence starts with the atoms located in one of the tweezers. Coherent splitting and recombining are achieved by an adiabatic change of the traps’ positions. We proposed two configurations for the splitters and combiners, utilizing either a two-tweezer or a three-tweezer setup.
@@
where $\sigma$ is the waist of the Gaussian beam creating each of the tweezers, $V_{0}$ is the potential depth, and $\Delta(t)$ is the relative depth difference (also referred to as detuning) between the two traps. Splitting of the wavefunction is achieved by changing the separation between the traps, $d(t)$, and the detuning in the following manner:
@@
The three-tweezer splitting scheme introduced in Ref. [2] is similar to the two-tweezer scheme but also allows the detection of erroneous performance. Within the scope of this paper, the differences between two- and three-tweezer splitters are not important, as the operation in both scenarios is described by the same unitary matrix.
@@
The guided interferometer architecture utilizing optical tweezers presents several benefits compared to free-fall light-pulse interferometry. In light-pulse interferometers, the atoms are in free fall, necessitating a larger apparatus for longer interrogation times. Conversely, in guided interferometers, the interrogation time is not dependent on the apparatus size, enabling a more compact system. Another limitation of light-pulse interferometry is that the atoms are in free-fall. In the proposed tweezer AIF, the atom’s position is entirely determined by the tweezer position, which allows for arbitrary trajectories during phase accumulation. Due to its enhanced sensitivity and arbitrary atom trajectory, the tweezer AIF can measure effects in classical and quantum gravity, e.g measuring the gravitational constant $G$ [45], testing the Newtonian law at small distances [45], measuring the gravitational redshift [39] and searching for evidence of the gravitational-field quantization [46, 47]. Most importantly, in the context of this paper, guided interferometers are sensitive to proper time differences between the arms [39], while existing light-pulse interferometers are inherently insensitive to proper time [41]. Moreover, an optical tweezer AIF working at a magic wavelength with an atomic clock is ideally suited to measure the gravitational time dilation, as we demonstrate in section III.
@@ III Tweezer clock interferometry
Optical tweezer AIFs, like those discussed in section II, can be used to measure phases arising from proper time differences if the atom has at least two internal states. This is due to the fact that an atom with a single internal degree of freedom cannot produce a periodic signal necessary for measuring proper time [30, 48]. Interferometers employing only a single internal state can only measure effects stemming from a difference of the gravitational potential at different location, analogous to a gravitational Aharonov-Bohm effect [49]. Such observations were made using atomic and neutron interferometry [50]. Atomic clock interferometry was previously proposed as a method to measure proper time differences [38, 42, 39, 48]. However, these proposals suffer from the same limitation of free falling AIF discussed in section II. Previous discussions of trapped AIFs [38, 39] were mostly concerned with calculating the redshift induced phase difference, without considering the questions of the splitting dynamics or the phase’s measurement protocol. In what follows, we present a proposal for a guided ACIF, using optical tweezers at a magic wavelength, including an analysis of its available obeservables. Our scheme inherits the advantages of the tweezer AIF, namely long probing duration and high position accuracy. Additionally, the measurement of the proper time difference is insensitive to relative intensity fluctuations, which makes it applicable to current tweezer technology.
@@
Our guided ACIF scheme is depicted in Fig. 1. It is a modification of the scheme discussed in section II, and can be applied with both the two- and three-tweezer approaches. Unlike the original tweezer AIF, it employs an atom which can be used as an atomic clock, i.e., with two long-lived internal states, denoted by |$\ket{g}$ and |$\ket{e}$. We choose the tweezers to operate at the magic wavelength of the clock, where both clock states experience the same trap potential. The atom is initially prepared in the ground state |$\ket{g}$. Before the wavepacket splitting stage, a $\pi/2$-pulse of an optical field, which is resonant with the |$\ket{g}\leftrightarrow\ket{e}$ transition, is applied. This pulse generates the superposition $\frac{1}{\sqrt{2}}(\ket{g}+\ket{e})$.
Following the $\pi/2$-pulse, the spatial splitting of the clock wavefunction occurs using either two or three tweezers. As detailed in section II, achieving balanced splitting relies on adiabatic following to reach the symmetric ground state of a symmetric double-well potential. However, in the modified scheme the splitting occurs in the direction defined by gravity. The presence of a gravitational potential disrupts this spatial symmetry, potentially leading to imbalanced splitting. Numerical simulations indicate that Earth’s gravitational potential results in a probability distribution of 44% and 56% for atoms being in the higher and lower arms, respectively, thus diminishing the visibility of interference fringes. To counteract this, applying a magnetic field gradient to negate the gravitational force during splitting and recombination is effective. Altering the depth of the tweezers during the operation can also compensate for the potential imbalance. Alternatively, conducting the splitting and recombining processes in a plane perpendicular to gravity and subsequently repositioning the arms at different heights can eliminate gravity’s influence. For the remainder of this analysis, we assume the imbalance in splitting is not significant and proceed with the evaluation using balanced atomic beam-splitters.
Each interferometer arm is moved to a different position in the gravitational field. Therefore, each part of the clock wavefunction experiences a different proper time. Before the recombining phase, a $\pi$-pulse on the |$\ket{g}$-|$\ket{e}$ transition is applied to only one of the arms. This pulse is necessary for the interferometer fringe visibility to scale linearly with the redshift, which is essential for making it detectable in a realistic experiment. Since the transitions between the clock states are driven by an optical field, it is straightforward to overlap the driving beam with only one of the tweezers to accomplish this step. The sequence ends with the spatial recombining of the wavefunction, as described in section II, followed by another $\pi/2$-pulse on the clock degrees of freedom to close the Ramsey sequence. The measured observables are the internal state of the atom and the spatial output port. As we demonstrate below, both quantities are required for a measurement of proper time differences in a coherent superposition of an atom.
@@
After the splitting and before the recombining, we define the energy difference $\Delta E$ (and the corresponding angular frequency $\omega_{0}$) between the two clock states in path 2 as
$\Delta E=\hbar\omega_{0}=E_{e;2}-E_{g;2}\,.$ (7)
@@
$E_{e;1}-E_{g;1}=\Delta E+\hbar\epsilon$ (8)
with $\epsilon=\frac{\omega_{0}}{c^{2}}\Delta\varphi$ being the angular frequency difference due to the gravitational redshift. Our goal for the ACIF sequence is to allow determination of $\epsilon$. It is important to note that in the way we wrote Eqs.(7-8) we explicitly assumed that the energy difference between the clock states does not depend on the depth of the confining optical potential, which is justified when the trap is operated at the magic wavelength.
@@
$U_{\text{phase}}(T)=\begin{pmatrix}1&0&0&0\\ 0&e^{-i\delta T}&0&0\\ 0&0&e^{i(\Delta+\epsilon)T}&0\\ 0&0&0&e^{i(\Delta-\delta)T}\end{pmatrix}\,,$ (12)
where we omit a global phase and define the detuning between the two lower state, $\delta=(E_{g;2}-E_{g;1})/\hbar$, and the drive detuning, $\Delta=\omega-\omega_{0}$, with $\omega$ being the frequency of the field driving the transitions involving the internal states. Eq. 12 is written in a rotating frame defined by the transformation
$U_{\text{rot}}(T)=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&e^{i\omega T}&0\\ 0&0&0&e^{i\omega T}\end{pmatrix}\,.$ (13)
@@
|$\ket{\psi_{f}}=U_{\text{rot}}^{\dagger}(T)U_{\pi/2}^{\dagger}U_{BS}^{\dagger}U_{\pi,l}U_{\text{phase}}(T)U_{BS}U_{\pi/2}\ket{g;1}\,.$ (14)
Note that we omit an initial $U_{\text{rot}}(0)$ since it is an identity.
In previous proposals of ACIF schemes, there is only one available observable for measurements of the interference pattern of the atomic phase due to proper time differences between the arms - either the spatial output port [31, 42, 38], or the internal clock state [39]. A unique feature of the proposed tweezer ACIF is that it has two spatial output ports and complete Ramsey sequence over the clock states, both exhibit coherent oscillations. The probability of the atoms to exit the interferometer from the upper port is given by
@@
$\begin{split}&P_{1}(\psi_{f})=P_{g;1}+P_{e;1}=|\langle g;1|\psi_{f}\rangle|^{2}+|\langle e;1|\psi_{f}\rangle|^{2}=\\ &=\frac{1}{2}\Biggl{[}1-\sin\Bigl{[}T\bigl{(}\frac{E_{g;2}-E_{g;1}}{\hbar}-\frac{\epsilon}{2}\bigr{)}\Bigr{]}\sin\Bigl{[}T\bigl{(}\Delta-\frac{\epsilon}{2}\bigr{)}\Bigr{]}\Biggr{]}\,.\end{split}$ (15)
@@
A crucial requirement for an interferometric measurement is the ability to verify that the atom was indeed in a coherent superposition during the sequence. The result of Eq.(15) allows this, since it exhibits coherent oscillations between the exit ports as a function of $T$. In contrast, if the wavepacket collapses randomly to one of the paths and the state becomes completely mixed, the exit port probability is $P_{1,\text{mixed}}(\psi_{f})=\frac{1}{2}$. This means that measuring oscillations around the probability of $\frac{1}{2}$ as a function of time verifies the coherence of the atomic wavefunction.
The primary drawback of measuring the spatial output port is its exit probability’s reliance on the difference between the eigenenergies: $E_{g;1}-E_{g;2}$. This dependency makes the observable vulnerable to relative intensity fluctuations of the trap beams. To address this issue, there are two strategies. Firstly, one could impose strict requirements on laser intensity stability throughout the interferometer’s duration to ensure minimal noise in this measurement. In conventional atomic interferometry, this is the sole approach since one needs to determine the relative phase between the paths, which is done by measuring the phase of the exit port oscillations.
In our case, however, we only need to verify the coherence of the split wavepacket, which allows for a simpler approach. Intensity differences between the tweezers (intentional or unintentional) can introduce essentially a random phase in the first sine term of Eq.(15). Given $x=\frac{1}{2}-A\sin(\phi)$ with $A\in[0,0.5]$ and the phase $\phi$ having a uniform random distribution in $[0,2\pi]$, the probability distribution of $x\in[0.5-A,0.5+A]$ is ${P(x)=\frac{1}{\pi\sqrt{A^{2}-(0.5-x)^{2}}}}$. In the case of Eq.(15), ${A=\frac{1}{2}\sin\Bigl{[}T\bigl{(}\Delta-\frac{\epsilon}{2}\bigr{)}\Bigr{]}\Biggr{]}}$ is set by the choice of $T$ and $\Delta$ while $\epsilon$ can be neglected. $T$ and $\Delta$ are well controlled in the experiment, and therefore $A$ can be regarded as constant. Thus, if the wavepacket is coherently split and maintained, the distribution of the estimator of $P_{1}$ should align with $P(x)$ across a sufficiently large dataset. Since $P_{1}$ changes from run to run, it is necessary to estimate it with sufficient accuracy in each run. This requires interferometeric runs with several atoms. As we show in the next section, 10 atoms per run are already enough to generate a clear distinction between coherent and non-coherent wavepacket.
With this approach we are still able to measure the gravitational redshift, $\epsilon$, thanks to to the availability of an additional observable - the probability of the atoms to be in the clock ground state at the end of the sequence (regardless of the output port). It is given by
@@
$\begin{split}P_{g}(\psi_{f})&=P_{g;1}+P_{g;2}=|\langle g;1|\psi_{f}\rangle|^{2}+|\langle g;2|\psi_{f}\rangle|^{2}\\ &=\frac{1}{2}\Bigg{[}1+\sin\Bigl{[}T\bigl{(}\Delta-\frac{\epsilon}{2}\bigr{)}\Bigr{]}\sin\Bigl{[}T\frac{\epsilon}{2}\Bigr{]}\Biggr{]}\,.\end{split}$ (16)
@@
As is readily seen from Eq.(16), in contrast to the case of measuring the spatial output port, the probability to finish in the ground state depends only on $\Delta$ and $\epsilon$, but not on the eigenenergies separately. Thanks to that, this observable is insensitive to relative fluctuations in the depths of the tweezers, which would shift the eigenenergies, but do not affect $\omega_{0}$ and $\Delta$, as long as we operate at the magic wavelength. This means that for measuring the internal clock state, the main noise source limiting guided AIFs is highly suppressed.
@@
The redshift is extracted from Eq.(16). By scanning the waiting time, $T$, the probability to find the atoms in the ground state oscillates due to interference of the two paths. The oscillation frequency of this interference pattern is set by the detuning of the driving field relative to the clock frequency, $\Delta$, shifted by $\frac{\epsilon}{2}$. Choosing ${\Delta\gg\frac{2\pi}{T},\epsilon}$ ensures that the oscillations can be observed by scanning the time from $T$ to $T+2\pi/\Delta$, during which the amplitude of the oscillations is almost constant and given by